Raymond Mining Corporation has 10.1 million shares of common stock outstanding, 450,000 shares of 5% $100 par value preferred stock outstanding, and 175,000 7.50% semiannual bonds outstanding, par value $1,000 each. The common stock currently sells for $49 per share and has a beta of 1.55, the preferred stock currently sells for $98 per share, and the bonds have 15 years to maturity and sell for 116% of par. The market risk premium is 8.9%, T-bills are yielding 4%, and Raymond Mining’s tax is 38%. The YTM before taxes is 5.880%. What is the WACC? (Enter your answers in whole dollars. Omit % sign in your response.)
Answer #1

rate positively ..

Rate of equity =
Using CAPM rate of equity = Risk free rate + market risk premium * beta
=4%+8.9%*1.55
17.795%
Rate of debt (after tax)
we have to use financial calculator to compute YTM
Put in calcualtor
PV -1160
FV 1000
PMT 1000*7.5%/2 37.5
N 30
Compute I 2.94%
YTM = 5.88%
tax rate = 38%
therefore rate of debt (after tax) = 3.65%
rate of debt (after tax) = 3.65%
rate of preferred stock = Annual dividend/Current price
=5/98
5.10%
Working for rest of question
Computation of Weight and WACC
Market value weight Cost of capital weight * cost
Source
equity 494900000 =10.1mil*49 66.70% 17.80% 11.87%
debt 203000000 =175000*1000*116% 27.36% 3.65% 1.00%
preferred stock 44100000 =450000*98 5.94% 5.10% 0.30%
742000000 13.17%
WACC= 13%
Ans=            13